OpenStudy (anonymous):

what is the inverse of x^3+10 ?

OpenStudy (anonymous):

so you have y = x^3 + 10, all you have to do is switch the variables x = y^3 + 10 and solve for y x-10 = y^3 y = cubed root of (x-10)

OpenStudy (anonymous):

is that the same as (x-10)^1/3 ?